# GaINeR: Geometry-Aware Implicit Neural Representation for Image Editing

### v1

@@ Abstract
Implicit Neural Representations (INRs) have become an essential tool for modeling continuous 2D images, enabling high-fidelity reconstruction, super-resolution, and compression. Popular architectures such as SIREN, WIRE, and FINER demonstrate the potential of INR for capturing fine-grained image details. However, traditional INRs often lack explicit geometric structure and have limited capabilities for local editing or integration with physical simulation, restricting their applicability in dynamic or interactive settings. To address these limitations, we propose GaINeR (Geometry-Aware Implicit Network Representation), a novel framework for 2D images that combines trainable Gaussian distributions with a neural network-based INR. For a given image coordinate, the model retrieves the K nearest Gaussians, aggregates distance-weighted embeddings, and predicts the RGB value via a neural network. This design enables continuous image representation, interpretable geometric structure, and flexible local editing, providing a foundation for physically aware and interactive image manipulation. The official implementation of our method is publicly available at https://github.com/WJakubowska/GaINeR.
Figure 1: GaINeR enables intuitive, geometry-aware image editing by manipulating Gaussian embeddings. By moving the Gaussians in space, our encoding–decoding pipeline produces semantically consistent and photometrically coherent transformations, demonstrating controllable image editing through learned Gaussian representations.
@@ 1 Introduction
Implicit Neural Representations, also known as coordinate-MLPs, have fundamentally reshaped how continuous signals are represented and processed across domains ranging from computer vision to computational physics [klocek2019hypernetwork, sitzmann2020implicit, takikawa2023neural]. Unlike traditional discrete representations tied to fixed spatial resolutions, INRs parameterize signals as continuous functions via neural networks, providing resolution-independent representations with high expressiveness and memory efficiency. While several architectures, such as SIREN [sitzmann2020implicit], GAUSS [ramasinghe2022beyond], WIRE [saragadam2023wire], FINER [liu2024finer], and positional encoding [tancik2020fourier] have improved the ability of INRs to capture high-frequency details in images, they typically lack explicit geometric structure, limiting the possibility of local editing and integration with physical simulations.
@@
Recent work, MiRaGe [waczynskamirage], addresses image editability by representing 2D images by a Gaussian Splatting-based model, enabling intuitive modifications and coupling with physics engines. Each Gaussian acts as a discrete component, allowing local transformations and perspective changes. However, because the representation is discrete, MiRaGe suffers from producing artifacts when the physical engine significantly alters the geometry of objects in a 2D image. The problem can be partially solved by training very small Gaussians, but it requires dedicated parameters that must be manually tuned. A Gaussian Splatting-based model cannot provide a truly continuous image representation.
@@
To overcome these limitations, we propose Geometry-Aware Implicit Network Representation (GaINeR), the first model that combines the continuous representation of INR for 2D images with trainable Gaussian embeddings (see Fig. 1). GaINeR enables editing 2D images and integration with physical simulation models, see Fig. 2. We achieve this by explicitly conditioning the implicit function on a geometric structure defined by a set of specific trainable Gaussian components, each parameterized by position, scale, and a unique learnable feature embedding. The core INR network then acts as a sophisticated decoder. Given a query coordinate $(x,y)$, the network retrieves and aggregates the features of the $k$ nearest Gaussians using a distance-weighted mechanism, see Fig. 3. The MLP then uses this aggregated geometric and feature information to synthesize the final, high-fidelity RGB colors. This design allows precise spatial manipulation of the image, as Gaussian positions and shapes can be adjusted to produce smooth, real-time deformations. Additionally, the Gaussians can act as discrete physical entities, enabling the entire 2D representation to interact naturally with standard physics engines for dynamic simulation and animation. As a result, GaINeR achieves state-of-the-art performance in 2D image reconstruction, while maintaining fully continuous and editable representations, see Fig. 4 and experimental section.
@@
  • We propose using Gaussians as feature providers conditioning an INR decoder, enabling high-fidelity and continuous color reconstruction.
  • We extend the INR framework by introducing a geometric component that allows direct and flexible 2D image modification and integrates with a physical simulation engine for interactive effects.
@@ 2 Related work
INR for images
INR architectures often remain simple, typically consisting of a single Multi-Layer Perceptron (MLP), with recent efforts focused on enhancing capacity through specialized embedding techniques [chen2022tensorf, muller2022instant], unique activation functions [sitzmann2020metasdf], refined rendering methods [barron2021mip], and regularization strategies [yang2023freenerf].
Implicit Neural Representations (INRs), or coordinate-based MLPs, have fundamentally transformed the representation and processing of continuous signals across various domains. The main direction of INRS’s development for 2D images is to adapt various activation functions and training procedures. SIREN [sitzmann2020implicit] uses sine activations alongside a frequency hyperparameter $\omega$, enabling smooth fitting and stable training. While effective for continuous signals, its fixed-frequency multiplier limits versatility, resulting in lower accuracy when modeling details across diverse frequencies [kaniafresh]. GAUSS [ramasinghe2022beyond] employs a Gaussian activation function providing smooth localized representations that are beneficial for signal reconstruction and denoising. However, its aperiodic nature hinders the effective modeling of sharp, high-frequency details. WIRE [saragadam2023wire] utilizes the Gabor wavelet activation, combining sinusoidal and Gaussian components to capture localized frequencies. Although effective for modeling high-frequency details within specific spatial regions, its complex-valued formulation increases computational demands, slowing both training and inference. FINER [liu2024finer] modifies SIREN with a frequency-variable activation, broadening the representable frequency spectrum while retaining SIREN’s benefits. However, FINER is sensitive to bias initialization and complicates the training process. Lastly, Positional Encoding (PE) [tancik2020fourier] maps inputs into a higher-dimensional space using sinusoidal functions to mitigate the spectral bias common in MLPs.
@@
Explicit and Editable Representations
Although INRs excel at compression and continuous representation, their inherent implicit nature severely limits direct interaction and editing. Conversely, explicit representations prioritize geometry and manipulability. $3D$ Gaussian Splatting (3DGS) [kerbl3Dgaussians] recently revolutionized novel view synthesis by employing a vast collection of explicit, $3D$ anisotropic Gaussians, achieving real-time rendering speeds. This approach naturally allows direct manipulation of Gaussian primitives, offering superior editability compared to traditional volumetric methods. Following this, GaussianImage [zhang2024gaussianimage] adapted the core GS concept to a $2D$ image representation, demonstrating competitive compression and quality. However, while these Gaussian-based methods introduce necessary explicit geometry, they still rely on simple color and opacity stored directly within each primitive. Consequently, they lose the continuous spatial coherence and superior color fidelity offered by a dedicated INR decoder. Models such as MiRaGe [waczynskamirage] attempted to address the editability challenge for $2D$ images by placing $2D$ Gaussian representations in a $3D$ space to enable physical interactions such as bending and folding.
@@ 3 Method
While GaINeR is introduced in the context of 2D image reconstruction and editing, its formulation is inherently dimension-agnostic. Since Gaussian embeddings and KNN-based aggregation operate in continuous Euclidean space, the same mechanism naturally extends to three-dimensional domains [muller2022instant]. Consequently, GaINeR generalizes seamlessly from 2D image representations to 3D volumetric modeling, as discussed in Section 3.5.
3.1 Classical Implicit Neural Representations
Implicit Neural Representations provide a continuous and compact way to encode signals such as images, shapes, or radiance fields using neural networks. Instead of storing discrete samples (e.g., pixels or voxels), an INR models the underlying signal as a function $f_{\theta}$ that directly maps input coordinates to output values. This formulation enables subpixel-level precision, smooth interpolation, and differentiability with respect to spatial coordinates.
@@ 3.1 Classical Implicit Neural Representations
Figure 4: Comparison of PSNR obtained on a Singapore image from DIV2K dataset. All models were trained for 30 k iterations, yet our method rapidly achieves the highest PSNR within the first 1–2 k iterations, often already surpassing all other approaches.
@@
3.2 GaINeR: Geometry-Aware Implicit Network Representation
Our proposed GaINeR extends the classical INR formulation by introducing a geometry-consistent embedding mechanism based on a family of learnable Gaussian distributions. Each 2D coordinate in the image domain is represented not as a fixed input vector, but as a continuous Gaussian embedding that encodes its local spatial context. This design enables geometry-aware reasoning and smooth interpolation while preserving the functional nature of implicit representations.
@@ 3.2 GaINeR: Geometry-Aware Implicit Network Representation
3.3 Training Objective
The training of GaINeR jointly optimizes the parameters of the Gaussian embedding $\{(\Sigma_{i},\mathbf{e}_{i})\}_{i=1}^{N}$, the multi-resolution hashgrid encoder $\mathcal{H}(\cdot)$, and the MLP decoder $f_{\theta}$ that reconstructs the target image.
Given a training set $\{(\mathbf{x}_{j},\mathbf{c}_{j})\}_{j=1}^{M}$, where $M\in N$ is a subset of Gausses in training dataset, $\mathbf{x}_{j}=(x_{j},y_{j})$ denotes pixel coordinates, and $\mathbf{c}_{j}=(r_{j},g_{j},b_{j})$ are the corresponding RGB values, the model is trained to minimize the Smooth L1 reconstruction loss:
@@ 3.3 Training Objective
Alpha-based Sampling and Mask Reconstruction.
For images containing an alpha channel, i.e., RGBA inputs, we utilize the alpha map as a spatial sampling mask during training to improve computational efficiency. Specifically, only pixels with nonzero alpha values are included in the training set $\{(\mathbf{x}_{j},\mathbf{c}_{j})\}_{j=1}^{M}$, which focuses optimization on visible regions and avoids redundant computation in transparent areas.
Figure 5: Overview of 3D reconstruction and rendering with GaINeR. Input images are first segmented using Segment Anything (SAM) [kirillov2023segment] and depth is estimated with Depth-Pro [bochkovskiy2025depthpro]. The model is trained in 2D, but during inference the Gaussian means $\mu_{i}$ are lifted into 3D using the estimated depth values. The decoder is extended to predict both color $\mathbf{c}$ and density $\sigma$, enabling volumetric rendering analogous to NeRF [mildenhall2021nerf]. This allows the model to synthesize consistent 3D views from novel camera perspectives, as illustrated at the bottom of the figure.
@@
3.4 Image Editing
During inference, the hashgrid encoder $\mathcal{H}$ is no longer required, since the learned Gaussian embeddings $\mathbf{e}_{i}$ already encode multi-scale spatial information acquired during training. Thus, the model operates directly on the optimized Gaussian parameters $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i},\mathbf{e}_{i})\}_{i=1}^{N}$, enabling efficient and interpretable scene editing.
@@ 3.4 Image Editing
3.5 2D To 3D Transformations
The proposed framework (Fig. 5) naturally generalizes beyond 2D image reconstruction. By extending the MLP decoder to additionally predict a density term $\sigma$,
@@
the model integrates with a volumetric rendering formulation analogous to NeRF [mildenhall2021nerf].
In this case, the rendering equation is expressed as
@@
Each Gaussian component is then lifted into three-dimensional space. The local embedding for any query position $\mathbf{x}$ is computed through the same Gaussian-weighted KNN aggregation $\mathbf{e}_{\mathbf{KNN}}(\mathbf{x},\mathcal{G}).$ As in the 2D formulation, shifting the Gaussian means $\boldsymbol{\mu}_{i}\mapsto\boldsymbol{\mu}_{i}^{\prime}$ directly modifies the interpolated embedding field $\mathbf{e}_{\mathbf{KNN}}(\mathbf{x},\mathcal{G})$, thereby altering the reconstructed volumetric scene while preserving photometric coherence.
Since the decoder $f_{\theta}$ depends solely on the aggregated embeddings rather than absolute coordinates, these transformations yield physically consistent and spatially smooth changes in appearance. Consequently, GaINeR provides a unified framework for continuous, differentiable editing across both 2D and 3D domains.
@@ 4 Experiments
We evaluated our proposed method across three tasks: image reconstruction, image editing, and 2D-to-3D lifting. The image editing task encompassed both user-driven transformations and physically simulated deformations. For reconstruction, we employed two standard benchmarks: Kodak [kodak_dataset] and DIV2K  [Agustsson_2017_CVPR_Workshops]. The Kodak dataset contains 24 natural images of resolution $768\times 512$, covering diverse scenes such as portraits, landscapes, and textures. The DIV2K validation set comprises 100 high-resolution images subjected to 2x bicubic downscaling, with spatial resolutions ranging from 408 x 1020 to 1020 x 1020, and is widely used for evaluating image restoration and super-resolution methods.
4.1 Quantitative comparisons
@@
Table 1: Quantitative comparison of our method with existing approaches on the Kodak and DIV2K datasets. Our method achieves the highest reconstruction quality across both datasets
Table  1 reports the average PSNR and MS-SSIM values obtained on both datasets. We compared our approach against WIRE [saragadam2023wire], SIREN [sitzmann2020implicit], I-NGP [muller2022instant], NeuRBF [chen2023neurbfneuralfieldsrepresentation], 3DGS  [kerbl3Dgaussians], GaussianImage  [zhang2024gaussianimage] and MiraGe [waczynskamirage]. In both cases, our method demonstrates a significant improvement in PSNR compared to previous approaches. Specifically, compared to the strongest existing method (MiraGe), our approach improves the average PSNR from 59.52 to 77.09 on the Kodak dataset, and from 54.54 to 62.20 on DIV2K, while maintaining near-perfect MS-SSIM values. This significant gain indicates that our representation captures image details with much higher fidelity and preserves structural consistency more effectively than prior implicit representations.
To compare performance, we trained six models (SIREN [sitzmann2020implicit], WIRE [saragadam2023wire], Gauss [ramasinghe2022beyond], FINER [liu2024finer], GaussianImage [zhang2024gaussianimage] and MiraGe [waczynskamirage]) for 30k iterations each to ensure a fair and consistent comparison. As shown in Fig. 4, our model surpasses all competing methods already from the initial iterations, achieving significantly higher PSNR even at the earliest stages of training. Remarkably, after only about 1–2k iterations (depending on the scene), our method reaches PSNR values higher than those obtained by other approaches after 30k iterations of training.
@@
Model Kodak dataset DIV2K dataset
@@
Table 2: Comparison of Gaussian-based methods evaluated using a unified configuration of 500k Gaussians (no densification) and 30k optimization iterations. Our method achieves the best performance.
4.2 Qualitative comparisons
@@ 4.3 Editing comparisons
We conducted a series of experiments on images from the DIV2K dataset, exploring various types of spatial modifications. For consistency across methods, we evaluate MiraGe using two camera configurations: one and mirror. The one camera setup corresponds to the same viewpoint configuration employed in our GaINeR model, ensuring direct comparability. The mirror camera provides a reflective viewpoint that captures mirrored scene geometry and appearance. As illustrated in Fig. 7, we evaluate the visual effects of bending, stretching, cutting and displacing Gaussian components. MiraGe exhibits severe distortions under these transformations, primarily due to its reliance on Gaussian primitives for direct rendering. In contrast, our GaINeR framework produces visually consistent and realistic results across all conditions, as it utilizes Gaussian components solely for encoding rather than for rendering.
4.4 Physics comparisons
@@
Figure 8: Qualitative comparison of physics-based animations. Elastic Flower: MiraGe’s render exhibits transparencies, while ours remains opaque. Elastic Beans: MiraGe produces spiky boundary artifacts, which are absent in our result. Sand Apple: As the object disintegrates, MiraGe’s rendering again shows structural artifacts, unlike GaINeR.
@@
To demonstrate applicability in dynamic scenarios, we treat the centers of our learned Gaussians as material points in a Material Point Method (MPM) simulation driven by the Taichi Elements [taichi_elements] engine. Fig. 8 compares our method against MiraGe, with full animations provided in the supplementary material. The results highlight the benefits of our continuous representation. The results show that MiraGe’s discrete representation produces geometric artifacts under deformation. In contrast, GaINeR renders coherent surfaces even under significant deformation. Since our MLP decoder is conditioned on the geometric structure, it can robustly interpolate visual information from sparse or distorted particle distributions, avoiding the visual discontinuities inherent to discrete representations.
4.5 2D-to-3D Lifting
@@
We performed a series of experiments demonstrating the capability of GaINeR to lift 2D representations into 3D space. As illustrated in Fig. 9, GaINeR reconstructs realistic and geometrically consistent 3D models by leveraging the spatially coherent nature of Gaussian embeddings and the KNN-based aggregation mechanism extended to three dimensions.
@@
We evaluated our method on various objects from the Tanks and Temples [knapitsch2017tanksandtemples], LLFF [mildenhall2019llff], and DTU [aanaes2012dtu] datasets (see also Fig. 5). Due to the continuity of the Gaussian representation in 3D, GaINeR also produces accurate depth maps for rendered views. For these experiments, depth supervision was provided using depth maps estimated by Depth-Pro [bochkovskiy2025depthpro], which were normalized to ensure consistent metric ranges across datasets.
@@
Figure 9: Results of 2D-to-3D lifting with GaINeR. Objects are rendered from multiple novel viewpoints (left to right), and the corresponding depth maps are shown in the top-left corners of each rendering. Some floaters appear due to the inconsistencies in generated depth maps by Depth-Pro.

### v2

# Equal contribution

@@ Abstract
Implicit Neural Representations (INRs) are widely used for modeling continuous 2D images, enabling high-fidelity reconstruction, super-resolution, and compression. Architectures such as SIREN, WIRE, and FINER demonstrate their ability to capture fine image details. However, conventional INRs lack explicit geometric structure, limiting local editing, and integration with physical simulation. To address these limitations, we propose GaINeR (Geometry-Aware Implicit Network Representation), a novel framework for 2D images that combines trainable Gaussian distributions with a neural network-based INR. For a given image coordinate, the model retrieves the K nearest Gaussians, aggregates distance-weighted embeddings, and predicts the RGB value via a neural network. This design enables continuous image representation, interpretable geometric structure, and flexible local editing, providing a foundation for physically aware and interactive image manipulation. Our method supports geometry-consistent transformations, seamless super-resolution, and integration with physics-based simulations. Moreover, the Gaussian representation allows lifting a single 2D image into a geometry-aware 3D representation, enabling depth-guided editing. Experiments demonstrate that GaINeR achieves state-of-the-art reconstruction quality while maintaining flexible and physically consistent image editing. The official implementation of our method is publicly available at https://github.com/WJakubowska/GaINeR.
Figure 1: Given a single image, GaINeR reconstructs a manipulable representation that enables geometry-aware editing. Our method supports object-level edits (top left), 2D to 3D transformations (top right), editing 3D reconstruction (bottom left), and upscaling (bottom right) while preserving semantic structure and photometric coherence.
@@ 1 Introduction
Implicit Neural Representations (INRs) have transformed the modeling of continuous signals across domains such as computer vision and computational physics [klocek2019hypernetwork, sitzmann2020implicit, takikawa2023neural]. Instead of discrete grids tied to fixed spatial resolutions, INRs represent signals as continuous functions parameterized by neural networks, enabling resolution-independent representations with high expressiveness and compact memory usage. While architectures such as SIREN [sitzmann2020implicit], GAUSS [ramasinghe2022beyond], WIRE [saragadam2023wire], FINER [liu2024finer], and positional encoding [tancik2020fourier] improve the ability of INRs to capture high-frequency image details, they typically lack explicit geometric structure, limiting local editing and integration with physical simulations.
@@
Recent work, MiRaGe [waczynskamirage], addresses image editability by representing 2D images by a Gaussian Splatting-based model, enabling intuitive modifications and coupling with physics engines. Each Gaussian acts as a discrete component, allowing local transformations and perspective changes. However, because the representation is discrete, MiRaGe suffers from producing artifacts when the physical engine significantly alters the geometry of objects in a 2D image. The problem can be partially solved by training very small Gaussians, but it requires dedicated parameters that must be manually tuned. Additionally a Gaussian Splatting-based model cannot provide a truly continuous image representation.
@@
To overcome these limitations, we propose Geometry-Aware Implicit Network Representation (GaINeR), the first model that combines the continuous representation of INR for 2D images with trainable Gaussian embeddings. GaINeR enables editing 2D images, and 2D to 3D manipulations (see Fig. 1, Fig. 2). We achieve this by explicitly conditioning the implicit function on a geometric structure defined by a set of specific trainable Gaussian components, each parameterized by position, scale, and a unique learnable feature embedding. The core INR network then acts as a decoder. Given a query coordinate $(x,y)$, the network retrieves and aggregates the features of the $k$ nearest Gaussians using a distance-weighted mechanism (see Fig. 3). The MLP then uses this aggregated geometric and feature information to synthesize the final, high-fidelity RGB colors. This design allows precise spatial manipulation of the image, as Gaussian positions and shapes can be adjusted to produce smooth, real-time deformations. Additionally, the Gaussians can act as discrete physical entities, enabling the entire 2D representation to interact naturally with standard physics engines for dynamic simulation and animation. As a result, GaINeR achieves state-of-the-art performance in 2D image reconstruction, while maintaining fully continuous and editable representations. Importantly, due to the locality of Gaussian conditioning, our representation naturally extends beyond the 2D image plane. By treating the image as a continuous surface embedded in 3D space, GaINeR enables lifting a single 2D image into a geometry-aware 3D representation. This allows depth-guided Gaussian displacement and 3D-consistent editing. Furthermore, as a fully continuous representation, GaINeR inherently supports arbitrary-resolution image synthesis, enabling seamless super-resolution and upscaling without retraining.
@@
  • We extend the INR framework by introducing a geometric component that allows direct and flexible 2D image modification and integrates with a physical simulation engine for interactive effects. Thanks to its fully continuous formulation, GaINeR also supports arbitrary-resolution image synthesis, enabling seamless super-resolution and upscaling without retraining.
  • We demonstrate that our geometry-aware formulation enables single-image 3D-aware transformations via depth-guided Gaussian displacement, allowing consistent spatial manipulation beyond the 2D image plane.
@@ 2 Related work
INR for images INR architectures often remain simple, typically consisting of a single Multi-Layer Perceptron (MLP), with recent efforts focused on enhancing capacity through specialized embedding techniques [chen2022tensorf, muller2022instant], unique activation functions [sitzmann2020metasdf], refined rendering methods [barron2021mip], and regularization strategies [yang2023freenerf].
Implicit Neural Representations (INRs), or coordinate-based MLPs, have fundamentally transformed the representation and processing of continuous signals across various domains. The main direction of INRs development for 2D images is to adapt various activation functions and training procedures. SIREN [sitzmann2020implicit] uses sine activations alongside a frequency hyperparameter $\omega$, enabling smooth fitting and stable training. While effective for continuous signals, its fixed-frequency multiplier limits versatility, resulting in lower accuracy when modeling details across diverse frequencies [kaniafresh]. GAUSS [ramasinghe2022beyond] employs a Gaussian activation function providing smooth localized representations that are beneficial for signal reconstruction and denoising. However, its aperiodic nature hinders the effective modeling of sharp, high-frequency details. WIRE [saragadam2023wire] utilizes the Gabor wavelet activation, combining sinusoidal and Gaussian components to capture localized frequencies. Although effective for modeling high-frequency details within specific spatial regions, its complex-valued formulation increases computational demands, slowing both training and inference. FINER [liu2024finer] modifies SIREN with a frequency-variable activation, broadening the representable frequency spectrum while retaining SIREN’s benefits. However, FINER is sensitive to bias initialization and complicates the training process. Lastly, Positional Encoding (PE) [tancik2020fourier] maps inputs into a higher-dimensional space using sinusoidal functions to mitigate the spectral bias common in MLPs.
@@
Explicit and Editable Representations Although INRs excel at compression and continuous representation, their inherent implicit nature severely limits direct interaction and editing. Conversely, explicit representations prioritize geometry and manipulability. $3D$ Gaussian Splatting (3DGS) [kerbl3Dgaussians] recently revolutionized novel view synthesis by employing a vast collection of explicit, $3D$ anisotropic Gaussians, achieving real-time rendering speeds. This approach naturally allows direct manipulation of Gaussian primitives, offering superior editability compared to traditional volumetric methods. Following this, GaussianImage [zhang2024gaussianimage] adapted the core GS concept to a $2D$ image representation, demonstrating competitive compression and quality. However, while these Gaussian-based methods introduce necessary explicit geometry, they still rely on simple color and opacity stored directly within each primitive. Consequently, they lose the continuous spatial coherence and superior color fidelity offered by a dedicated INR decoder. Models such as MiRaGe [waczynskamirage] attempted to address the editability challenge for $2D$ images by placing $2D$ Gaussian representations in a $3D$ space to enable physical interactions such as bending and folding.
@@ 3 Method
While GaINeR is introduced in the context of 2D image reconstruction and editing, its formulation is inherently dimension-agnostic. Since Gaussian embeddings and KNN-based aggregation operate in continuous space, the same mechanism naturally extends to three-dimensional domains [muller2022instant]. Consequently, GaINeR generalizes seamlessly from 2D image representations to 3D volumetric modeling.
Classical Implicit Neural Representations Implicit Neural Representations provide a continuous and compact way to encode signals such as images, shapes, or radiance fields using neural networks. Instead of storing discrete samples (e.g., pixels or voxels), an INR models the underlying signal as a function $f_{\theta}$ that directly maps input coordinates to output values. This formulation enables subpixel-level precision, smooth interpolation, and differentiability with respect to spatial coordinates.
@@
Geometry-Aware Implicit Network Representation Our method extends the classical INR formulation by introducing a geometry-consistent embedding mechanism based on a family of learnable Gaussian distributions. Each 2D coordinate in the image domain is represented not as a fixed input vector, but as a continuous Gaussian embedding that encodes its local spatial context. This design enables geometry-aware reasoning and smooth interpolation while preserving the functional nature of implicit representations.
@@
Figure 4: Results of 2D-to-3D lifting with GaINeR. We elevate flat 2D images into fully volumetric 3D representations using monocular depth maps. The final column shows the reconstructed geometry rendered from a novel viewpoint. Our method generalizes well across complex paintings and masked objects.
Training Objective The training of GaINeR jointly optimizes the parameters of the Gaussian embedding $\{(\Sigma_{i},\mathbf{e}_{i})\}_{i=1}^{N}$, the multi-resolution hashgrid encoder $\mathcal{H}(\cdot)$, and the MLP decoder $f_{\theta}$ that reconstructs the target image.
Given a training set $\{(\mathbf{x}_{j},\mathbf{c}_{j})\}_{j=1}^{M}$, where $M\in N$ is a subset of Gaussians in training dataset, $\mathbf{x}_{j}=(x_{j},y_{j})$ denotes pixel coordinates, and $\mathbf{c}_{j}=(r_{j},g_{j},b_{j})$ are the corresponding RGB values, the model is trained to minimize the Smooth L1 reconstruction loss:
@@
Alpha-based Sampling and Mask Reconstruction. For images containing an alpha channel, i.e., RGBA inputs, we utilize the alpha map as a spatial sampling mask during training to improve computational efficiency. Specifically, only pixels with nonzero alpha values are included in the training set $\{(\mathbf{x}_{j},\mathbf{c}_{j})\}_{j=1}^{M}$, which focuses optimization on visible regions and avoids redundant computation in transparent areas.
@@
Image Editing During inference, the hashgrid encoder $\mathcal{H}$ is no longer required, since the learned Gaussian embeddings $\mathbf{e}_{i}$ already encode multi-scale spatial information acquired during training. Thus, the model operates directly on the optimized Gaussian parameters $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i},\mathbf{e}_{i})\}_{i=1}^{N}$, enabling efficient and interpretable scene editing.
@@
Figure 5: Overview of 3D reconstruction with GaINeR. Input images are segmented with SAM [kirillov2023segment] and lifted to 3D using depth estimated by Depth-Pro [bochkovskiy2025depthpro]. The Gaussian means are projected into 3D while the decoder predicts color $c$ and density $\sigma$, enabling NeRF-style volumetric rendering and novel-view synthesis, and direct 3D scene modifications.
2D To 3D Transformations The proposed framework (Fig. 5) naturally generalizes beyond 2D image reconstruction. By extending the MLP decoder to additionally predict a density term $\sigma$,
@@
the model integrates with a volumetric rendering formulation analogous to NeRF [mildenhall2021nerf]. In this case, the rendering equation is expressed as
@@
Each Gaussian component is then lifted into three-dimensional space. The local embedding for any query position $\mathbf{x}$ is computed through the Gaussian-weighted KNN aggregation $\mathbf{e}_{\mathbf{KNN}}(\mathbf{x},\mathcal{G}).$ However the KNN utilizes Mahalanobis distance instead of Euclidean distance for the search for nearest neighbors. As in the 2D formulation, shifting the Gaussian means $\boldsymbol{\mu}_{i}\mapsto\boldsymbol{\mu}_{i}^{\prime}$ directly modifies the interpolated embedding field $\mathbf{e}_{\mathbf{KNN}}(\mathbf{x},\mathcal{G})$, thereby altering the reconstructed volumetric scene while preserving photometric coherence.
Since the decoder $f_{\theta}$ depends solely on the aggregated embeddings rather than absolute coordinates, these transformations yield physically consistent and spatially smooth changes in appearance. Consequently, GaINeR provides a unified framework for continuous, differentiable editing across both 2D and 3D domains, enabling 2D-to-3D lifting followed by direct 3D scene editing.
@@ 4 Experiments
We evaluated our proposed method across three tasks: image reconstruction, image editing, and 2D-to-3D lifting. The image editing task encompassed both user-driven transformations and physically simulated deformations. For reconstruction, we employed two standard benchmarks: Kodak [kodak_dataset] and DIV2K  [Agustsson_2017_CVPR_Workshops]. Additional ablation studies, runtime analysis, and implementation details are provided in the supplementary material. The Kodak dataset contains 24 natural images of resolution $768\times 512$, covering diverse scenes such as portraits, landscapes, and textures. The DIV2K validation set comprises 100 high-resolution images subjected to 2x bicubic downscaling, with spatial resolutions ranging from 408 x 1020 to 1020 x 1020, and is widely used for evaluating image restoration and super-resolution methods.
@@
Table 1: Quantitative comparison on the Kodak and DIV2K datasets. Our method achieves the highest reconstruction quality across both datasets in terms of PSNR and MS-SSIM.
Quantitative and Qualitative Evaluation Table  1 reports the average PSNR and MS-SSIM values obtained on both datasets. We compared our approach against WIRE [saragadam2023wire], SIREN [sitzmann2020implicit], I-NGP [muller2022instant], NeuRBF [chen2023neurbfneuralfieldsrepresentation], 3DGS  [kerbl3Dgaussians], GaussianImage  [zhang2024gaussianimage] and MiraGe [waczynskamirage]. In both cases, our method demonstrates a significant improvement in PSNR compared to previous approaches. Specifically, compared to the strongest existing method (MiraGe), our approach improves the average PSNR from 59.52 to 77.09 on the Kodak dataset, and from 54.54 to 62.20 on DIV2K, while maintaining near-perfect MS-SSIM values. This significant gain indicates that our representation captures image details with much higher fidelity and preserves structural consistency more effectively than prior implicit representations.
@@
Model Kodak DIV2K
@@
Table 2: Comparison with Gaussian-based approaches on the Kodak and DIV2K datasets. Our method achieves the highest reconstruction quality in terms of PSNR and MS-SSIM.
@@
Editing comparisons We conducted a series of experiments on images from the DIV2K dataset, exploring various types of spatial modifications. For consistency across methods, we evaluate MiraGe using two camera configurations: one and mirror. The one camera setup corresponds to the same viewpoint configuration employed in our GaINeR model, ensuring direct comparability. The mirror camera provides a reflective viewpoint that captures mirrored scene geometry and appearance. As illustrated in Fig. 7, we evaluate the visual effects of bending, stretching, cutting and displacing Gaussian components. MiraGe exhibits severe distortions under these transformations, primarily due to its reliance on Gaussian primitives for direct rendering. In contrast, our GaINeR framework produces visually consistent and realistic results across all conditions, as it utilizes Gaussian components solely for encoding rather than for rendering.
@@
Figure 8: Qualitative comparison of physics-based animations. Elastic Flower: MiraGe’s render exhibits transparencies, while ours remains opaque. Elastic Beans: MiraGe produces spiky boundary artifacts, which are absent in our result. Sand Apple: As the object collapses, MiraGe’s rendering shows structural artifacts, unlike GaINeR.
@@
Physics comparisons To demonstrate applicability in dynamic scenarios, we treat the centers of our learned Gaussians as material points in a Material Point Method (MPM) simulation driven by the Taichi Elements [taichi_elements] engine. Fig. 8 compares our method against MiraGe, with full animations provided in the supplementary material. While MiraGe’s discrete representation produces geometric artifacts and spiky boundaries under stress, GaINeR renders coherent surfaces even under significant deformation. Since our MLP decoder is conditioned on the geometric structure, it can robustly interpolate visual information from sparse or distorted particle distributions, avoiding the visual discontinuities inherent to discrete representations. We further showcase this robustness in Fig. 9, where GaINeR successfully handles complex, multi-material interactions, such as fluid mechanics and granular dynamics, maintaining continuous, high-fidelity textures despite extreme topological changes over time.
Figure 9: Physical simulation of complex multi-material interactions using GaINeR. By treating learned Gaussian centers as material points in a Material Point Method engine, we simulate granular dynamics (sand) and fluid mechanics (water) simultaneously. Our continuous MLP decoder maintains cohesive surfaces, fluid-like blending, and high-frequency textures even under severe topological deformations over time.
@@
2D-to-3D Lifting We performed a series of experiments demonstrating the capability of GaINeR to lift 2D representations into 3D space. As illustrated in Fig. 4, GaINeR reconstructs realistic and geometrically consistent 3D models by leveraging the spatially coherent nature of Gaussian embeddings and the KNN-based aggregation mechanism extended to three dimensions.
@@
We evaluated our method on various objects from the LLFF [mildenhall2019llff] and DTU [aanaes2012dtu] datasets (see also Fig. 5). Due to the continuity of the Gaussian representation in 3D, GaINeR also produces accurate depth maps for rendered views. For these experiments, depth supervision was provided using depth maps estimated by Depth-Pro [bochkovskiy2025depthpro], which were normalized to ensure consistent metric ranges across datasets.
@@
Upscaling
Figure 10: Visual comparison of GaINeR with 3D Gaussian Splatting and MiraGe at ×8, ×10, and ×12 scaling factors. Our method maintains higher perceptual quality, whereas competing approaches introduce Gaussian artifacts.
Image upscaling aims to reconstruct high-resolution images from low-resolution inputs while preserving fine details and structural consistency, a task that becomes particularly challenging at high magnification factors where artifacts and over-smoothing often appear. Fig. 10 shows a qualitative comparison of GaINeR with 3D Gaussian Splatting and MiraGe at ×8, ×10, and ×12 scales. As an implicit neural representation (INR), GaINeR generates high-resolution outputs without relying on fixed discrete grids. As illustrated in the figure, it preserves fine textures and sharp edges even at extreme magnifications, while 3D Gaussian Splatting and MiraGe exhibit visible degradation artifacts, including Gaussian-like patterns and over-smoothed regions. All experiments were conducted on the DIV2K dataset, where the input images were obtained by bicubic downscaling of the original images by a factor of 8.